Hello,

I need help solving and writing an equation for this problem:

A number is decreased by 45. The result is then divided by 12. Then 20 is added to this new result to give a final result of 25. What is the number?

Thank you

Answer by stanbon(75887)

A number is decreased by 45. The result is then divided by 12. Then 20 is added to this new result to give a final result of 25. What is the number?
---------
Let the number be "x".
---------------
EQUATION:
[(x-45)/12]+20 = 25
(x-45)/12 = 5
x-45 = 60
x = 105
=============
Cheers,
Stan H.